### v1

# AnyBipe: An End-to-End Framework for Training and Deploying Bipedal Robots Guided by Large Language Models

Abstract
Training and deploying reinforcement learning (RL) policies for robots, especially in accomplishing specific tasks, presents substantial challenges. Recent advancements have explored diverse reward function designs, training techniques, simulation-to-reality (sim-to-real) transfers, and performance analysis methodologies, yet these still require significant human intervention. This paper introduces an end-to-end framework for training and deploying RL policies, guided by Large Language Models (LLMs), and evaluates its effectiveness on bipedal robots. The framework consists of three interconnected modules: an LLM-guided reward function design module, an RL training module leveraging prior work, and a sim-to-real homomorphic evaluation module. This design significantly reduces the need for human input by utilizing only essential simulation and deployment platforms, with the option to incorporate human-engineered strategies and historical data. We detail the construction of these modules, their advantages over traditional approaches, and demonstrate the framework’s capability to autonomously develop and refine controlling strategies for bipedal robot locomotion, showcasing its potential to operate independently of human intervention.
I Introduction
With the integration of advanced algorithms, enhanced physical simulations, and improved computational power, robotics has made significant strides[1, 2]. These innovations enable robots to perform tasks ranging from industrial automation to personal assistance with unprecedented efficiency and autonomy[3, 4]. As industrial robotics matures, researches focus on humanoid robots, particularly in replicating human-like characteristics and enabling robots to perform human-oriented tasks[5]. Correspondingly, bipedal robots can be used to simulate the morphology of human lower limbs, thus providing a method to explore the locomotion skills of humanoid robots [6, 7].
Control strategies for bipedal robots typically leverage either traditional control methods or reinforcement learning (RL) methods[8, 9]. Traditional approaches rely on problem abstraction, modeling, and detailed planning, while RL employs reward functions to iteratively guide robots toward task completion. Through repeated interactions with the environment, RL enables robots to refine control strategies and acquire essential skills, particularly excelling in trial-and-error learning in simulated environments, where robots adapt to complex terrains and disturbances.
Despite these advancements, training and deploying RL algorithms remains challenging. Effective reward design requires careful consideration of task-specific goals and the incorporation of safety constraints for real-world applications[10, 11]. This complexity demands significant engineering effort in training, testing, and iterative refinement. Although reward shaping and safe RL offer potential solutions[12, 13], they often rely on prior experience, complicating the reward design process. Furthermore, bridging the gap between simulations and real-world conditions—the ”Sim-to-Real” challenge—remains difficult[14]. Techniques such as domain randomization[15], which randomizes physical parameters to enhance agent robustness, and observation design, which facilitates task transfers across varied terrains, prove formidable, but still require real-world testing and human feedback for model selection and fine-tuning[7, 16, 17].
Hohe integration of large language models (LLMs) into robotics offers the prospect of reducing human work in such process. Known for their capabilities in code generation and missing planning, LLMs are being widely applied to complex robotics applications[18]. For instance, they play a pivotal role in embodied intelligence by enabling the dynamic creation of action tasks[19]. Recent developments have explored the utility of LLMs in improving reward function design and refining from simulation performance—key areas that reduce the need for human intervention[20]. However, a comprehensive framework that designs suitable policies and automatically implements properly trained models in real-world remains lacking. To address this issue, we propose a novel framework that leverages LLMs to optimize the entire training-to-deployment process. The framework minimizes human involvement, enables autonomous design, training, and deployment of RL algorithms, and supports exploring strategies from scratch or enhancing existing ones.
We present AnyBipe111https://github.com/sjtu-mvasl-robotics/AnyBipe.git, the first end-to-end framework for training, simulating, and deploying bipedal robots, guided by LLMs. This framework consists of three interconnected modules and embodies a closed-loop process for continual refinement and deployment. Our main contribution are listed as follows:
1.
Minimized Human Intervention: The proposed framework operates autonomously, requiring only minimal initial user input to oversee the entire process from training to deployment. No additional human intervention is needed during the workflow.
2.
LLM-Guided Reward Function Design: Leveraging large language models, the framework generates suitable reward functions from predefined prompts. Through iterative refinement, it allows users to design customized RL reward functions from the ground up.
3.
Incorporation of Prior Knowledge in Training: The framework enables the integration of pre-trained models and conventional control algorithms, which enhances RL training stability and facilitates the migration of traditional control implementations into the proposed system.
4.
Real-World Feedback Integration via Homomorphic Evaluation: This module converts real-world sensor and actuator feedback into formats compatible with simulation, enabling LLMs to bridge the gap between the training environment and real-world deployment. As a result, it allows for adaptive adjustments to the reward function based on actual feedback.
Experimental trials on bipedal robots traversing both flat and complex terrains have shown that AnyBipe can significantly enhance real-world deployment performance. Compared to manually designed reward functions, those generated by AnyBipe lead to faster convergence and more stable training outcomes, also preventing falling into high-reward but low-usability solutions. Furthermore, homomorphic real-world feedback allows LLMs to understand the impact of the reward function on actual deployment performance, thereby improving the corresponding design and improving Sim-to-Real performance.
This paper is organized as follows: First, we introduce the design principles and implementation methods of the framework modules in detail. Then, we describe the experiments used to verify the role of each module and express the training and deployment effects of the framework. Finally, we summarize our contributions and propose future research directions for our framework.
II Related Works
Reinforcement Learning for Bipedal Robots. Reinforcement learning (RL) has achieved significant success in enabling legged robots to learn locomotion and motion control through data-driven methods, allowing them to adapt to diverse environmental challenges[21, 22, 23, 24, 25, 26]. Research has introduced RL-based locomotion strategies, training in simulation environments like MuJoCo[27] and Isaac Gym[28]. Recent studies have shifted these techniques, originally designed for quadruped robots and robot arms, to bipedal robots like Cassie[29, 30]. Additional approaches explore imitation learning[31], motion planning[32], and robust RL strategies[7], enabling bipedal robots to perform tasks like running, stair climbing, and complex maneuvers[33, 34, 35]. Building on these advancements, our work utilizes the Isaac Gym environment and proposes a semi-supervised RL framework to transfer the original work and effectively improve the learning performance.
Large Language Model Guided Robotics. Large language models (LLMs) have demonstrated considerable capabilities in task understanding[36, 37], semantic planning[38, 39], and code generation[40, 41, 42], allowing them to be effectively integrated into a variety of robotic tasks. LLMs automate environmental analysis[43, 44], reward functions design[45, 46], and task-action mappint[47, 48]. However, challenges such as data scarcity, real-time performance, and real-world integration remain[49]. In addition, due to the lack of awareness of actual data, LLM-driven code design sometimes needs to be optimized through human feedback. Our framework addresses the existing limitations by adopting environmental characteristics and safety constraints as priors, and uniquely combines homomorphic feedback from real-world applications, reducing the need for human feedback in the process.
Sim-to-real Training and Deploying Techniques. The gap between simulated environments and real-world conditions, known as the “reality gap,” presents significant challenges for deploying RL strategies in robotics[50, 51]. Techniques such as domain randomization[52, 53, 54] and system identification[55, 56, 57] are widely used to address this issue. Researchers have proposed sim-to-real solutions for bipedal robots to handle tasks such as turning and walking[58, 59]. Recent work has also integrated LLMs to enhance environmental modeling and reward function design, making simulations more reflective of real-world complexity[20]. However, most approaches still rely on separate training in simulation and real-world evaluation, often using human feedback to assess sim-to-real effectiveness. Our work extends these techniques by introducing an evaluation loop that continuously monitors sim-to-real performance during deployment.
III Methods
In this section, we introduce the AnyBipe framework in detail. The framework consists of three modules aiming to automate reward design, training and simulation, and deployment feedback to minimize the need for manual work. The framework requires the robot’s URDF model, a basic task description, an RL training platform, and a ROS-based system for communication and control [60], along with an SDK for sensor and actuator data, as compulsory input. Optional elements include a manual reward function template, a teacher model realized with previous work, and custom environmental observations for evaluation. AnyBipe generates reward functions accordingly and trains policies supervised by the teacher model. A group of top strategies, determined by evaluation criteria and simulation results, are deployed via ROS. After validating in the ROS Gazebo simulation, policies that pass the safety check undergo real-world testing, and the best strategy is then selected. Evaluation result across various environments and the selected best policy are compiled to feedback, guiding iterative improvements in reward generation. The procedural steps are encapsulated in Algorithm 1.
III-A Module 1: LLM Guided Reward Function Design
We refer to the reward function design in the Eureka algorithm [45], which enables LLMs to autonomously improve and iterate reward functions with predefined environmental and task descriptions . However, initial usability issues require multiple iterations for viable training code. Furthermore, Eureka often overlooks discrepancies between training and real environments , resulting in computationally expensive but minimally effective reward functions that may induce undefined behaviors [61]. The framework also lacks comprehensive safety considerations for tasks such as bipedal movement, despite attempts to integrate safety through Reward-Aware Physical Priors (RAPP) and LLM-led Domain Randomization [20].
To address key issues, we developed a robust context establishment mechanism that effectively tackles underdesigned reward functions and safety constraints. Our approach classifies prompts into two categories: General and Task-Specific. For general tasks, we provide coding tips, function templates, and predefined templates that facilitate code compilation, training feedback, and evalu feedback. We also integrate reference tables from Isaac Gym for environmental variables like motor torque, torque limits, and feet height. These tables prevent LLMs from the customization of non-existent observations while enhancing the utilization of environmental variables in in order to ensure that LLMs consider actionable constraints during reward function design. Our experimental results affirm that LLMs can seamlessly incorporate these safety restrictions and environmental variables, thus designing highly effective reward functions, evidencing their exceptional context tracking and directive following capabilities [62, 63].
Furthermore, the Task-Specific module allows users to define custom prompts for specific tasks, facilitating the rapid generation of viable code and standardization of reward calculations. Users have the flexibility to use trainable artificial reward functions as templates or employ various computational paradigms to enhance the accuracy and applicability of reward assessments.
To refine the evaluation of designed reward functions for completing the improvement loop, we introduced a comprehensive reward function evaluation scheme. This scheme not only tracks changes in rewards and observations throughout the training but also integrates a homomorphic evaluation model to closely assess real-world robot performance. This model ensures a high correlation between real-world outcomes and theoretical reward functions, enabling LLMs to intuitively identify the most impactful components of the reward functions. Details on this model are expounded in subsection III-C, showcasing our commitment to precise and practical feedback mechanisms that enhance real-world applicability.s more significantly to real-world performance.
III-B Module 2: RL Training Adopting Reference Policy as Teacher
In this section, we detail the adaptations applied to guide RL training towards desired actions using frameworks provided by Legged Gym222https://github.com/leggedrobotics/legged_gym, employing the Proximal Policy Optimization (PPO) algorithm[64] as our foundation. We assume the existence of a baseline policy , which could be derived from traditional control methods or previous RL techniques.
To enhance the PPO algorithm, we modify the objective function given below:
where denotes the probability ratio, is an estimator of the advantage at time , is a small positive number, and is a coefficient that measures the divergence between the reference policy and PPO policy.
This integration enables control over the similarity between the trained policy and the reference policy. However, abstracting the reference policy into the same probabilistic framework as the PPO policy presents challenges. Despite this, given the deterministic nature of actions for a specific state and previous action , and assuming sufficient environmental observations, we can approximate the distribution as a Dirac distribution, and the differences between and can be described as follows:
The approximation uses the integral properties of the Dirac function and expression of KL divergence, which is omitted due to space constraints. With adequate observations, this approximation becomes reliable. This improvement can effectively prevent reinforcement learning from falling into a degradation trap or a local optimal solution.
In the AnyBipe framework, we have established a template for deploying existing policies as teacher functions, demonstrating integration of models deployed using PyTorch, ONNX, and traditional control algorithms implemented in C++ into our framework. This will allow users to transfer previous work to current work, or introduce pre-trained policies for simple tasks to accelerate the convergence of policies on complex tasks.
III-C Module 3: Deployment, Evaluation, and Feedback
In this section, we discuss optimal strategy selection, state estimator implementations, and methods to align evaluation with reward functions, leading to refined strategy generation by LLMs.
To shield our data from ineffective reward functions, we define essential training metrics—like command compliance and survival duration—as basic measurements , filtering samples that meet these benchmarks. For specific tasks such as bipedal stair climbing or standing, customization of observational metrics and coefficients is enabled in configuration files. We then select the top 15% of strategies for real-world deployment, prioritizing those that maximize combined metrics:
Strategies are deployed in a ROS-based robotic framework () using ONNX. Given that the training () and real environments () are isomorphic, we define a homomorphism , ensuring the real-world evaluation metric mirrors the reward function. Our automated script aligns reward structures with observed real-world data, informing users of any mismatches.
We also develop a basic state estimator for comprehensive robotic assessments, including metrics like step width and leg height. Models are first validated in Gazebo to confirm operational stability, with safety checks aligned with safety regulations proposed in subsection III-A. Successful models undergo real-world testing, with optimal models selected based on:
AnyBipe enables an autonomous cycle from reward function generation to training, deployment, and optimization, facilitating user-driven training of bipedal robot RL algorithms without pre-existing strategies.
IV Experiments
Our experiments were conducted on a six-degree-of-freedom bipedal robot from Limx Dynamics, shown in Fig. 4. GPT-4o was selected as the LLM base, with training performed on an NVIDIA GTX 3090Ti. We explored the robot’s locomotion on both flat and complex terrains over multiple experimental rounds. This section highlights key experiments on individual modules and overall training effectiveness. Due to space limitations, only critical results are presented here, with detailed data and charts available on our GitHub page.
Table I outlines key reward functions that maintain the same structure across both human-engineered and AnyBipe-generated rewards, differing only in scale, allowing direct performance comparison.
| Reward Name | Expression Form |
| Survival | |
| Tracking Linear Velocity | |
| Tracking Angular Velocity | |
| Success |
IV-A Module Analysis
We evaluated each module by testing prompt design and various augmentations using gpt-3.5-turbo, gpt-4, and gpt-4o with samples per batch. Success rates—defined as the proportion of samples processed successfully in the first iteration—were compared across four configurations: basic prompts, prompts with code references, safety regulations, and both. Fig. 5 shows that adding code references and safety regulations improves LLM performance in generating usable reward functions. As gpt-4 and gpt-4o show similar performance with complete prompts, gpt-4o is recommended as the primary model.
We validated the effect of incorporating safety regulations into reward functions by comparing models with and without safety prompts in Isaac Gym and Gazebo environments. Fig. 6 shows the operational postures and IMU spatial angles of safe models (c)(d) versus unsafe models (a)(b). Results indicate that, even without real-world feedback, safety prompts effectively constrain robot behavior, supporting practical deployment.
We evaluated teacher-guided models under artificial reward functions, training each for 5,000 iterations on complex terrain. The original model used only an artificial reward function, while the teacher-guided model was instructed by an operational ONNX model with . Performance was measured using reward success and terrain level. Results showed that the teacher-guided model had more stable training, with faster and less volatile reward growth.
For the Homomorphic evaluation part, the following table shows some evaluation indicators before and after the conversion, and the evaluation results under the best model. It can objectively reflect some differences from simulation to reality. Each column seperately represents reward function name, reward in isaac gym, homomorphic measurement in gazebo, in reality, and the mapped tracking result in reality (30 seconds of tracking).
| Name | Gym | Gazebo | Reality | Mapping (real/targ) |
| Track lin vel | 56.92 | 21.93 | 20.41 | 0.86 (1.0) m/s |
| Track ang vel | 36.73 | 15.60 | 14.43 | 0.06 (0.10) rad/s |
| Feet distance | -0.31 | -0.00 | -0.00 | m (m) |
| Standing still1 | -50.35 | -6.16 | - 11.20 | 5.8 (30) s |
| Survival time | 0.86 | 0.30 | 0.30 | 30 (30) s |
Note: 1Standing still requires robot to maintain speed 0 for certain time.
IV-B Framework Analysis
The experimental setup involved first establishing a baseline using a manually designed reward function. We then trained our model from scratch in flat and complex terrain environments, sed to simulate the general situation of users first using the AnyBipe framework. The robot was trained to track speed commands, avoid falls, and walk on various terrains using a simplified reward function with basic components like velocity tracking, action rate, and balance judgment, without scale-related information. Over five rounds ( samples each), the best model from each round was tested in Gazebo and real-world scenarios. The basic training framework occupies 7G of GPU memory, and the corresponding training time is 79 hours. As shown in Fig. 8, initial iterations had issues like excessive movement (Iteration 0) and unnatural joint postures (Iterations 1 and 2), but these were corrected by Iteration 5, achieving a natural gait.
To demonstrate the effectiveness of the LLM’s reward function improvements during training, we evaluate its performance using reward success and terrain level in tasks like velocity tracking, safe walking, and navigating complex terrains. After comparing five iterations with manually designed reward functions in Fig. 9, we found that the LLM’s reward function outperformed the manual version after just two iterations. Each subsequent iteration further enhanced training speed and performance, all achieved without human intervention.
To verify that the final trained policy not only performs well in the lab but also adapts to real-world terrains, we conducted walking tests across five different surfaces: the experimental site, carpet, hard ground, grass, and stairs. The experiments demonstrated that the model trained and deployed by AnyBipe possesses the ability to walk on various terrains. In contrast, the manually designed reward function used in the experiment failed to achieve Sim-to-real transfer, indicating that AnyBipe can independently resolve the Sim-to-real problem.
V Conclusion
AnyBipe proposes an end-to-end framework for training and deploying bipedal robots, which utilizes a state-of-the-art LLM to design reward functions for specific tasks. The framework provides interfaces that allow users to supply coding references and integrate pre-existing models to assist in training. Additionally, it incorporates feedback from both simulated and real-world test results, enabling the execution of Sim-to-Real tasks entirely without human supervision. We validated the effectiveness of each module, as well as the system’s ability to guide the robot in learning locomotion in complex environments, continuously improving the model by either designing new reward functions from scratch or refining existing ones. Furthermore, this framework exhibits potential for transferability to other robotic task planning scenarios. Our future work will focus on improving the current framework in three key areas: first, extending it to a broader range of robotic applications to verify its generalizability; second, testing its effectiveness across more tasks beyond locomotion; and third, enhancing the model evaluation process by incorporating visual and conceptional feedback to achieve more comprehensive state estimation.
APPENDIX
V-A Proof of section 3.2
We have KL divergence between normal distribution with PDF , and Dirac distribution with PDF , the KL divergence can be written as
We first prove that . Let , , where is the step function, we have
Therefore we can approximately say that
Then we show why we don’t calculate the normal KL divergence . This is because the term
is , there is no way for normal calculation. Since the backward propagation functions normally, we choose the inverse form, which approximately measures the difference between two distributions.
Also, in practical, we tend to ignore the term which controls the size of , using this term directly instead: